# Performance of a Superconducting Quantum Battery

### v1

@@ III Results and discussion
Figure 4(b) shows the instantaneous power $\mathcal{P}$, which continues to exhibit similar oscillatory behavior as observed in Figs. 2(b) and 3(b). The increased amplitude and frequency of oscillation with larger $\xi_{c}$ parallel the trends noted in ergotropy. This suggests that the power output of the system also improves, providing more energy within a given time frame. The correlation between ergotropy and instantaneous power reinforces the observation that an enhanced coupling energy directly translates into a greater efficiency of the quantum battery.
Figure 4(c) illustrates the capacity $\mathcal{K}$, which remains constant across the various values of $\tau$, as in Figs. 2(c) and 3(c). However, the value of $\mathcal{K}$ is directly influenced by $\xi_{c}$: increasing $\xi_{c}$ leads to a higher constant value of capacity throughout the domain of $\tau$. This suggests that while energy extraction fluctuates with ergotropy and power, the system’s information transmission capabilities are enhanced by larger $\xi_{c}$, maintaining a stable and robust performance across the domain of $\tau$.
Figure 4(d) examines the $l_{1}$-norm of coherence $\mathcal{C}_{l_{1}}$. Similar to the behavior observed in Figs. 2(d) and 3(d), $\mathcal{C}_{l_{1}}$ shows oscillatory behavior as $\xi_{c}$ increases. However, unlike ergotropy and instantaneous power, the amplitude of the oscillations in coherence slightly declines with increasing $\xi_{c}$, despite the frequency of oscillation increasing. This decline in amplitude suggests that although the qubits become more strongly coupled, leading to more frequent state transitions (i.e., more maxima), the overall ability of the system to maintain a high level of coherence diminishes slightly. Physically, this could reflect the increased interaction between the qubits causing more rapid fluctuations in coherence but limiting the maximum achievable coherence at any given time.
@@
Figure 5(a) depicts the plots of ergotropy $\mathcal{E}$ with respect to $\tau$. While the overall tendencies are similar to those in Figure 4(a), we observe that the maximum values of ergotropy are significantly lower, which means that the maximum amount of energy stored in the quantum battery is reduced. This is due to the smaller value of $\xi_{2}$, which represents the Josephson energy of the second qubit. With $\xi_{1}=1.5$ and $\xi_{2}=0.5$, the asymmetry in the Josephson energies of the qubits reduces the system’s ability to extract work. The oscillatory nature persists, with both the amplitude and frequency increasing as $\xi_{c}$ increases. However, the reduction in $\xi_{2}$ limits the overall ergotropy, as the second qubit contributes less energy to the system.
@@
Figure 5(b) shows the instantaneous power $\mathcal{P}$. Similar to the trends observed in ergotropy, the maximum values of $\mathcal{P}$ are lower compared to Fig. 4(b). This again reflects the reduced Josephson energy of the second qubit, which lowers the power output of the system. The oscillations in power increase in both amplitude and frequency as $\xi_{c}$ increases, consistent with the behavior of ergotropy. The coupling strength $\xi_{c}$ enhances the rate of energy transfer between the qubits, but the total available energy is constrained by the smaller value of $\xi_{2}$.
@@
Figure 5(c) illustrates the capacity $\mathcal{K}$, which also shows lower overall values. The reduced capacity is directly related to the reduced Josephson energy of the second qubit, $\xi_{2}=0.5$, which lowers the system’s ability to store and transfer information. As in Fig. 4(c), capacity increases with the quantity $\xi_{c}$, indicating that $\mathcal{K}$ is affected by both the coupling strength and the Josephson energies of the qubits.
@@
Figure 5(d) examines the $l_{1}$-norm of coherence $\mathcal{C}_{l_{1}}$. The general behavior of coherence is similar to that in Fig. 4(d), with oscillations increasing in frequency as $\xi_{c}$ increases. The amplitude of the oscillations in $\mathcal{C}_{l_{1}}$ is also declining as we increase the values of $\xi_{c}$. Moreover, while the coupling strength enhances the frequency of coherence oscillations, the asymmetry between the qubits’ Josephson energies ($\xi_{1}=1.5$ and $\xi_{2}=0.5$) reduces the maximum achievable coherence as compared with Fig. 4(d).
@@
The results in Fig. 5 highlight the impact of the asymmetry in Josephson energies between the two qubits on the system’s performance. With $\xi_{1}=1.5$ and $\xi_{2}=0.5$, the overall ergotropy, instantaneous power, and capacity are significantly reduced compared to the results in Fig. 4. This suggests that the energy imbalance between the qubits limits their collective ability to extract work, generate power, and store information, influencing the overall performance of the quantum battery. However, the coupling strength $\xi_{c}$ still plays a key role in increasing the frequency of oscillations in ergotropy, power, and coherence, demonstrating that stronger coupling facilitates more dynamic interactions between the qubits.
IV Discussion and Conclusions
@@
In this paper, we investigated the performance of SQB charging using a quantum system composed of two superconducting coupled qubits, specifically focusing on the variation of key thermodynamic and informational quantities: ergotropy $\mathcal{E}$, instantaneous power $\mathcal{P}$, capacity $\mathcal{K}$, and the $l_{1}$-norm of coherence $\mathcal{C}_{l_{1}}$. By systematically varying the coupling strength $\xi_{c}$ as well as the Josephson energies of the qubits $\xi_{1}$ and $\xi_{2}$, we have explored how these parameters influence the performance of the quantum battery.
@@
Our analysis demonstrates several significant findings. First, we observe that for all values of the coupling strength $\xi_{c}$, the system exhibits oscillatory behavior in ergotropy, instantaneous power, and coherence. These oscillations reflect the underlying quantum energy transfer dynamics between the two qubits. As $\xi_{c}$ increases, both the amplitude and frequency of these oscillations increase. This tendency means that stronger coupling enhances the energy exchange between the qubits and results in more rapid transitions within the system. However, in cases where there is an asymmetry between the Josephson energies of the qubits, particularly when $\xi_{2}$ is significantly smaller than $\xi_{1}$, the maximum achievable ergotropy and power are limited. This implies that a balance in the Josephson energies of the qubits is crucial for maximizing the system’s performance, particularly for work extraction and power generation.
@@ Appendix A Elements of thermal state
$\mathcal{R}_{th_{13}}=\mathcal{R}_{th_{31}}=\frac{\frac{B_{-}\left(\xi_{1}-\xi_{2}\right)}{2\alpha_{-}}+\frac{B_{+}\left(\xi_{1}+\xi_{2}\right)}{2\alpha_{+}}}{2\left(A_{-}+A_{+}\right)},$ (19)
@@
$\mathcal{R}_{th_{43}}=\mathcal{R}_{th_{34}}=\frac{\frac{B_{-}\left(\xi_{2}-\xi_{1}\right)}{2\alpha_{-}}+\frac{B_{+}\left(\xi_{1}+\xi_{2}\right)}{2\alpha_{+}}}{2\left(A_{-}+A_{+}\right)},$ (23)
with
@@
and
$\alpha_{\pm}=\sqrt{4\xi_{c}^{2}+(\xi_{1}\pm\xi_{2})^{2}},$
$A_{\pm}=\cosh\left(\frac{\alpha_{\pm}}{2T}\right),$
$B_{\pm}=\sinh\left(\frac{\alpha_{\pm}}{2T}\right).$

### v2

@@ III Results and discussion
Fig. 4(b) shows the instantaneous power $\mathcal{P}$, which continues to exhibit similar oscillatory behavior as observed in Figs. 2(b) and 3(b). The increased amplitude and frequency of oscillation with larger $\xi_{c}$ parallel the trends noted in ergotropy. This suggests that the power output of the system also improves, providing more energy within a given time frame. The correlation between ergotropy and instantaneous power reinforces the observation that an enhanced coupling energy directly translates into a greater efficiency of the SQB.
Fig. 4(c) illustrates the capacity $\mathcal{K}$, which remains constant across the various values of $\tau$, as in Figs. 2(c) and 3(c). However, the value of $\mathcal{K}$ is directly influenced by $\xi_{c}$: increasing $\xi_{c}$ leads to a higher constant value of capacity throughout the domain of $\tau$. This suggests that while energy extraction fluctuates with ergotropy and power, optimal performance is achieved with larger $\xi_{c}$, ensuring stable and robust performance across the range of $\tau$.
Fig. 4(d) examines the $l_{1}$-norm of coherence $\mathcal{C}_{l_{1}}$. Similar to the behavior observed in Figs. 2(d) and 3(d), $\mathcal{C}_{l_{1}}$ shows oscillatory behavior as $\xi_{c}$ increases. However, unlike ergotropy and instantaneous power, the amplitude of the oscillations in coherence slightly declines with increasing $\xi_{c}$, despite the frequency of oscillation increasing. This decline in amplitude suggests that although the qubits become more strongly coupled, leading to more frequent state transitions (i.e., more maxima), the overall ability of the system to maintain a high level of coherence diminishes slightly. Physically, this could reflect the increased interaction between the qubits causing more rapid fluctuations in coherence but limiting the maximum achievable coherence at any given time.
@@
Fig. 5(a) depicts the plots of ergotropy $\mathcal{E}$ with respect to $\tau$. While the overall tendencies are similar to those in Fig. 4(a), we observe that the maximum values of ergotropy are significantly lower, which means that the maximum amount of energy stored in the SQB is reduced. This is due to the smaller value of $\xi_{2}$, which represents the Josephson energy of the second qubit. With $\xi_{1}=1.5$ and $\xi_{2}=0.5$, the asymmetry in the Josephson energies of the qubits reduces the system’s ability to extract work. The oscillatory nature persists, with both the amplitude and frequency increasing as $\xi_{c}$ increases. However, the reduction in $\xi_{2}$ limits the overall ergotropy, as the second qubit contributes less energy to the system.
@@
Fig. 5(b) shows the instantaneous power $\mathcal{P}$. Similar to the trends observed in ergotropy, the maximum values of $\mathcal{P}$ are lower compared to Fig. 4(b). This again reflects the reduced Josephson energy of the second qubit, which lowers the power output of the system. The oscillations in power increase in both amplitude and frequency as $\xi_{c}$ increases, consistent with the behavior of ergotropy. The coupling strength $\xi_{c}$ enhances the rate of energy transfer between the qubits, but the total available energy is constrained by the smaller value of $\xi_{2}$.
@@
Fig. 5(c) illustrates the capacity $\mathcal{K}$, which also shows lower overall values. The SQB capacity is hardly influenced by the reduced Josephson energy of the second qubit, $\xi_{2}=0.5$, which lowers the system’s ability to store and transfer energy. In addition, the capacity is more sensitive to coupling strength $\xi_{c}$ in Fig. 4(c), the increase of $\xi_{c}$ will increase the SQB capacity. It means that we could enhance the capacity by increasing the coupling strength and the Josephson energies of the qubits.
@@
Fig. 5(d) examines the $l_{1}$-norm of coherence, denoted as $\mathcal{C}_{l_{1}}$. The general behavior of coherence is similar to that in Fig. 4(d), with oscillations increasing in frequency as $\xi_{c}$ rises. Additionally, the amplitude of the oscillations in $\mathcal{C}_{l_{1}}$ declines with increasing values of $\xi_{c}$. While the coupling strength enhances the frequency of coherence oscillations, the asymmetry between the qubits’ Josephson energies ($\xi_{1}=1.5$ and $\xi_{2}=0.5$) leads to a reduction in the maximum achievable coherence compared to Fig. 4(d). However, there is also a clear difference to be noticed. The symmetrical Josephson energies yield significantly higher coherence.
@@
The results in Fig. 5 highlight the impact of asymmetry in the Josephson energies of the two qubits on the system’s performance. With $\xi_{1}=1.5$ and $\xi_{2}=0.5$, the overall ergotropy, instantaneous power, and capacity are significantly reduced compared to the results in Fig. 4. This observation confirms that the energy imbalance between the qubits limits their collective ability to extract work, generate power, and store information, thereby influencing the overall performance of the SQB. The coupling strength parameter $\xi_{c}$ is beneficial as it increases the frequency of oscillations in ergotropy, power, and coherence, demonstrating that stronger coupling leads to more dynamic interactions between the qubits.
IV Scalability of SQBs
Scalability is a crucial aspect of practical quantum battery implementations, particularly in superconducting circuit architectures. While our analysis focuses on SQB as a fundamental building block, the principles established in this work can be extended to multi-qubit configurations.
One promising method for scalability is the utilization of collective charging protocols in multi-qubit superconducting systems. Previous studies have shown that collective effects, such as quantum synchronization and cooperative energy transfer, can lead to enhanced charging power and improved energy storage efficiency in quantum batteries [20, 59]. In a superconducting circuit, such collective charging could be realized through coherent coupling between multiple transmon qubits or by employing superconducting resonators as mediators for global interactions [60].
Another approach involves leveraging modular circuit architectures that integrate superconducting qubit arrays into larger networked configurations. Recent advances in superconducting quantum computing have demonstrated the feasibility of scaling up qubit arrays with high coherence times and tunable interactions, which could be exploited for enhanced quantum battery performance [61, 62]. By optimizing the connectivity between qubits and tuning Josephson energies appropriately, it may be possible to sustain coherent energy transfer across a larger quantum battery system while mitigating decoherence effects.
Additionally, flux-tunable qubits and superconducting metamaterials provide further scalability options by enabling dynamic control over interactions, allowing for adaptive charging protocols that optimize energy storage efficiency as the system scales [14, 63].
Future investigations should focus on the robustness of these multi-qubit architectures under realistic noise conditions and their potential advantages over classical and semi-classical charging mechanisms. Understanding the interplay between coherence, entanglement, and charging efficiency in large-scale superconducting quantum batteries will be essential for advancing their practical implementation.
V Experimental feasibility and optimization strategies
Our study is grounded in realistic superconducting circuit parameters, ensuring that the proposed SQB model aligns with current experimental capabilities in superconducting qubit technology. Recent advancements in superconducting quantum circuits have demonstrated precise tunability of coupling strengths and energy levels, enabling controlled interactions between qubits [61, 63]. Moreover, the analysis considers the tunability of Josephson energies and coupling strengths, both of which can be dynamically controlled in state-of-the-art superconducting platforms using external magnetic flux and circuit design optimizations [64, 65]. This flexibility ensures that our model can be implemented using existing fabrication techniques. Interestingly, the oscillatory behavior of ergotropy and power observed in our model is directly linked to experimentally accessible coherence times and coupling strengths in superconducting qubits. These characteristics have been extensively studied in transmon-based superconducting circuits, which exhibit coherence times on the order of tens to hundreds of microseconds, allowing sufficient time for energy transfer and extraction [50, 14].
Regarding the optimization strategies, we demonstrate that tuning the coupling strength $\xi_{c}$ enhances the frequency and amplitude of energy oscillations, offering a method to optimize power output in practical applications. This is consistent with previous studies showing that strong coupling enhances quantum energy transfer efficiency in superconducting systems [43]. Our results indicate that an optimal balance between the Josephson energies $\xi_{1}$ and $\xi_{2}$ is crucial for maximizing energy storage and extraction. Imbalances in Josephson energies create asymmetries in the system’s energy levels, which can limit the maximum extractable work. Similar effects have been observed in other superconducting qubit-based quantum thermodynamic models [66]. By analyzing the interplay between coherence, energy transfer, and information capacity, we offer insights into designing SQBs with improved performance, particularly in mitigating losses due to parameter asymmetry. Our findings align with prior research on coherence-driven quantum energy storage, reinforcing the role of quantum coherence as a resource for energy efficiency [11, 56].
Thus, this paper provides a theoretical foundation for engineering SQBs with controlled energy dynamics, paving the way for experimental validation and further optimization in quantum energy storage systems. With continued advancements in superconducting qubit fabrication and control, experimental demonstrations of SQBs based on coupled qubits may soon be realized, delivering optimal performance [20, 15, 67].
VI Discussion and Conclusions
@@
In this paper, we investigated the performance of SQB charging using a quantum system composed of two superconducting coupled qubits, specifically focusing on the variation of ergotropy $\mathcal{E}$, instantaneous power $\mathcal{P}$, capacity $\mathcal{K}$, and the $l_{1}$-norm of coherence $\mathcal{C}_{l_{1}}$. By systematically varying the coupling strength $\xi_{c}$ as well as the Josephson energies of the qubits $\xi_{1}$ and $\xi_{2}$, we have explored how these parameters influence the performance of the quantum battery.
@@
Our analysis demonstrates several significant findings. First, we observe that for all values of the coupling strength $\xi_{c}$, the system exhibits oscillatory behavior in ergotropy, instantaneous power, and coherence. These oscillations reflect the underlying quantum energy transfer dynamics between the two qubits. As $\xi_{c}$ increases, both the amplitude and frequency of these oscillations increase. This tendency means that stronger coupling enhances the energy exchange between the qubits and results in more rapid transitions within the system. The large coupling strength $\xi_{c}$ corresponds to stronger coherent interactions between the two qubits, resulting in higher-frequency oscillations in ergotropy, power, and coherence during the charging process. Efficient energy transfer can be adjusted to a significant value by selecting the coupling strength $\xi_{c}$ appropriately. However, in cases where there is an asymmetry between the Josephson energies of the qubits, particularly when $\xi_{2}$ is significantly smaller than $\xi_{1}$, the maximum achievable ergotropy and power are limited. This is because a significant disparity in $\xi_{1}$ and $\xi_{2}$ affects the energy gaps of the system, limiting the extent to which energy can be cyclically stored and extracted. The observed reduction in peak ergotropy and power is thus not a suppression of energy exchange per se but rather a constraint on how effectively energy can be harnessed for work extraction due to the system’s altered energy landscape. This implies that the Josephson energies of the qubits must be equal to ensure maximum energy transfer.
@@ Appendix A Elements of thermal state
$\mathcal{R}_{th_{13}}=\mathcal{R}_{th_{31}}=\mathcal{R}_{th_{24}}=\mathcal{R}_{th_{42}}=\frac{\frac{B_{-}\left(\xi_{1}-\xi_{2}\right)}{2\alpha_{-}}+\frac{B_{+}\left(\xi_{1}+\xi_{2}\right)}{2\alpha_{+}}}{2\left(A_{-}+A_{+}\right)},$ (19)
@@
$\mathcal{R}_{th_{34}}=\mathcal{R}_{th_{43}}=\frac{\frac{B_{-}\left(\xi_{2}-\xi_{1}\right)}{2\alpha_{-}}+\frac{B_{+}\left(\xi_{1}+\xi_{2}\right)}{2\alpha_{+}}}{2\left(A_{-}+A_{+}\right)},$ (23)
where the partition function is given by
@@
with $\alpha_{\pm}=\sqrt{4\xi_{c}^{2}+(\xi_{1}\pm\xi_{2})^{2}},$ $A_{\pm}=\cosh\left(\frac{\alpha_{\pm}}{2T}\right),$ and $B_{\pm}=\sinh\left(\frac{\alpha_{\pm}}{2T}\right).$
@@ Appendix C Closed-form expressions for performance metrics of SQB
Appendix D Parameter analysis
In this appendix, we provide a detailed discussion of the physical meaning and selection basis of the parameters used in our simulations to ensure clarity and reproducibility.
D.1 Selection of system parameters
The model considered in this paper involves several key parameters: $\xi_{1}$, $\xi_{2}$, $\xi_{c}$, and $T$, which play crucial roles in defining the system’s behavior.
Interaction parameters $\xi_{1}$ and $\xi_{2}$.
The parameters $\xi_{1}$ and $\xi_{2}$ represent the Josephson energies of the qubits in the quantum battery system, directly influencing energy transfer and coherence dynamics. Their selection is based on typical values observed in superconducting qubit setups, ensuring that the studied regime is experimentally feasible. Specifically:
  • In Fig. 2, $\xi_{1}$ is fixed at $1.5$ while $\xi_{2}$ is varied over the values 0.1, 0.5, 1, and 2. This range allows us to explore different regimes, from weak ($\xi_{2}=0.1$) to strong ($\xi_{2}=2$) Josephson energy. At lower values, the qubit exhibits slower charging dynamics and weaker coherence effects, whereas at higher values, it contributes more significantly to energy storage and coherence preservation.
  • In Fig. 3, $\xi_{2}$ is fixed at 1.5 while $\xi_{1}$ is varied over 0.1, 0.5, 1, and 2. This complementary analysis highlights the role of $\xi_{1}$ in the system’s performance. Similar to $\xi_{2}$, increasing $\xi_{1}$ enhances energy storage efficiency and coherence retention, while lower values correspond to weaker qubit contributions to the charging process.
  • The choice of $\xi_{1},\xi_{2}=1.5$ as a reference point, is motivated by previous theoretical studies [68, 69, 17, 70], where similar values were found to optimize and amplify the quantum correlations and quantum coherence.
Coupling strength $\xi_{c}$.
The parameter $\xi_{c}$ represents the coupling strength between the battery and its external driving mechanism. It is crucial in determining the rate of energy transfer.
  • In Fig. 4, we analyze the impact of varying $\xi_{c}$ while keeping $\xi_{1}=\xi_{2}=1.5$ fixed, allowing us to isolate its effect on energy dynamics. Specifically, we consider $\xi_{c}=0.1,0.5,1$, and $2$, covering a range from weak to strong coupling, which enables us to assess how increasing interaction strength influences coherence and power transfer.
  • In Fig. 5, we repeat this analysis with a different parameter set ($\xi_{1}=1.5,\xi_{2}=0.5$) to confirm robustness across different regimes. Again, we use $\xi_{c}=0.1,0.5,1$, and $2$, ensuring a broad exploration of interaction effects.
  • The choice of $\xi_{c}=0.05$ or $\xi_{c}=0.5$ in fixed cases (Figs. 2 and 3) aligns with typical weak and moderate coupling regimes.
Temperature $T$.
The temperature $T$ affects the thermal fluctuations within the system. Our simulations explore two representative cases:
  • $T=0.5$ in Fig. 2, representing a higher temperature scenario where thermal effects might influence performance.
  • $T=0.1$ in Figs. 35, representing a lower temperature regime, consistent with dilution refrigerator environments used in superconducting qubits.
D.2 Reproducibility considerations
The parameter values were chosen to reflect realistic conditions for quantum battery implementations, ensuring that the numerical results are both theoretically grounded and experimentally feasible. To enhance reproducibility, we explicitly state the parameter values in all figure captions. Our analysis focuses on comparing distinct scenarios, such as weak vs. moderate coupling or low vs. high temperature, to illustrate the influence of key parameters on quantum battery performance. This approach allows for a clearer interpretation of how different system parameters affect energy storage and transfer.